OpenStudy (ashleyy_smith):

~~I'll give a medal :))~~ Find an equation for the line that passes through the points, (3,−2) and , (−3,6).

OpenStudy (ashleyy_smith):

@FortyTheRapper

OpenStudy (nubeer):

do you know how to find a gradient?

OpenStudy (ashleyy_smith):

no :( @nubeer

OpenStudy (nubeer):

\[m=\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\] "m" is the gradient here

OpenStudy (ashleyy_smith):

oh okay. How would I plug in the numbers

OpenStudy (nubeer):

(3,-2) -- > these are x1 and y1 x1=3,y1=-2 (-3,6) x2=-3,y2=6

OpenStudy (nubeer):

i believe you can find "m" now

OpenStudy (ashleyy_smith):

I got - 8/6

OpenStudy (ashleyy_smith):

@nubeer

OpenStudy (nubeer):

gud. now to find any equation of line the formula is as following: \[y=mx + c\]

OpenStudy (nubeer):

you have found m right?

OpenStudy (ashleyy_smith):

Oh that's the answer? :O

OpenStudy (nubeer):

no :3

OpenStudy (nubeer):

i am jut trying to explain you step by step

OpenStudy (ashleyy_smith):

oh y = mx + c

OpenStudy (nubeer):

m = -8/6 (as you found above) y=mx+c now plug m here

OpenStudy (ashleyy_smith):

I dont get it :s

OpenStudy (nubeer):

:/

OpenStudy (nubeer):

which part u don't get?

OpenStudy (ashleyy_smith):

what I have to plug in y = mx + c

OpenStudy (nubeer):

the value "m" you just found above

OpenStudy (nubeer):

m is also known as gradient

OpenStudy (ashleyy_smith):

do I have to do something with that to get mx??

OpenStudy (nubeer):

-_- let me start over again

OpenStudy (nubeer):

you have two coordinates given to you in question.. i asked you to find the gradient. you got -8/6 so far correct?

OpenStudy (ashleyy_smith):

yes

OpenStudy (nubeer):

okay.. this was only the half solution.. so the next step is to know the formula to find the equation of line (which is asked in your question) Basically there are two formulas to find the equation of line \[y=mx+c\] OR \[y-y _{1}=m(x-x _{1})\] Are you familiar or studied any one of them at school or anywhere?

OpenStudy (ashleyy_smith):

no not yet :s

OpenStudy (nubeer):

okay.. well you can remember it now that you have to use any one of the formula above. so which formula you would like to choose?

OpenStudy (ashleyy_smith):

which one is easier? :o

OpenStudy (nubeer):

both are.. but for starters u can use y=mx+c

OpenStudy (ashleyy_smith):

oh okay

OpenStudy (nubeer):

y=mx + c here we usually have to find "m" and "c" we did find "m" what was it?

OpenStudy (ashleyy_smith):

-8/6

OpenStudy (nubeer):

okay now we plug this value in y=mx+c y = (-8/6)x +c

OpenStudy (nubeer):

now only the missing variable is c pick any coordinate given in your question

OpenStudy (ashleyy_smith):

ummm 3, -2

OpenStudy (nubeer):

okay.. put them in the equation now.. y= (-8/6)x +c find c here

OpenStudy (ashleyy_smith):

y= (-8/6)3 + -2?

OpenStudy (nubeer):

:3 no x =3, y=-2 pluggin in equation -2 = (-8/6) (3) +c -2 = -4 +c c =2 so y= (-8/6)x +2 this is your answer

OpenStudy (ashleyy_smith):

oh thankk uu :)